# Randomized block-Krylov subspace methods for low-rank approximation of matrix functions

### v1

@@ 4 Numerical experiments
In this section we compare the Krylov aware low-rank approximation (Algorithm 4) and Algorithm 2 (assuming exact matvecs with $f(\bm{A})$) and Algorithm 3 (inexact matvecs with $f(\bm{A})$). All experiments have been performed in MATLAB (version 2020a) and scripts to reproduce the figures are available at https://github.com/davpersson/Krylov_aware_LRA.git.
@@ 4.1.1 Exponential integrator
The following example is taken from [32]. Consider the following parabolic differential equation
$\begin{split}&u_{t}=\kappa\Delta u+\lambda u\text{ in }[0,1]^{2}\times[0,2]\\ &u(\cdot,0)=\theta\text{ in }[0,1]^{2}\\ &u=0\text{ on }\Gamma_{1}\\ &\frac{\partial u}{\partial\bm{n}}=0\text{ on }\Gamma_{2}\end{split}$
@@
Hence, we can use Algorithm 4 to compute $\bm{Q}_{s}$ and $\bm{T}_{q}$ and use them to efficiently construct a rank $60$ approximation to $\exp(t\bm{A})$ for any $t$ with almost no additional cost.
@@ 4.1.2 Estrada index
For an (undirected) graph with adjacency matrix $\bm{A}$ the Estrada index is defined as $\operatorname{tr}(\exp(\bm{A}))$. It is used to measure the degree of protein folding [12]. One can estimate the Estrada index of a network by the Hutch++ algorithm or its variations [6, 10, 28, 31], which requires computing a low-rank approximation of $\exp(\bm{A})$. Motivated by the numerical experiments in [28] we let $\bm{A}$ be the adjacency matrix of Roget’s Thesaurus semantic graph [1].
@@
Figure 1: Comparing relative error (19) for the the approximations returned by Algorithm 4 without truncation (untruncated), Algorithm 4 with truncation back to rank $k$ (truncated), Algorithm 3, and Algorithm 2. The black line shows the optimal rank $k$ approximation relative Frobenius norm error. The rank parameter $k$ is visible as titles in the figures. In all experiments we set $\ell=k$.
@@ 4.1.3 Quantum spin system
We use an example from [10, Section 4.3], a similar example is found in [6], in which we want to approximate $\exp(-\beta\bm{A})$ where
@@
Estimating the partition function $Z(\beta)=\operatorname{tr}(\exp(-\beta\bm{A}))$ is an important task in quantum mechanics [35], which once again can benefit from computing a low-rank approximation of $\exp(-\beta\bm{A})$.
@@ 4.1.4 Synthetic example for the matrix logarithm
Figure 2: Comparing relative error (19) for the the approximations returned by Algorithm 4 without truncation (untruncated), Algorithm 4 with truncation back to rank $k$ (truncated), Algorithm 3, and Algorithm 2. The black line shows the optimal rank $k$ approximation relative Frobenius norm error. The rank parameter $k$ is visible as titles in the figures. In all experiments we set $\ell=k+5$.
4.2 Comparing relative errors
In this section we compare the error returned by Algorithm 4, Algorithm 2, and Algorithm 3. If $\bm{C}$ is a low-rank approximation returned by one of the algorithms then we compare the relative error
@@ 4.2 Comparing relative errors
In all experiments we set the parameters in Algorithm 4 and Algorithm 3 to be $\ell=k$ or $\ell=k+5$ and $s=r$ so that the total number of matrix vector products with $\bm{A}$ is $2\ell s$. When we run Algorithm 3 we compute matvecs with $f(\bm{A})$ exactly, which cannot be done in practice. Hence, the results from this algorithm are only used as a reference for Algorithm 4 and Algorithm 3. The results are presented in Figure 1 for $\ell=k$ and Figure 2 for $\ell=k+5$. All results confirm that Algorithm 4 returns a more accurate approximation than Algorithm 3, and can even be more accurate than the Algorithm 2. Note that for the example given in Section 4.1.3 the error for the untruncated version of the approximation returned by Algorithm 4 stagnates. This is because the error from the approximation of the quadratic form dominates the error. In this case, $r$ should be chosen larger than $s$. However, we leave it as an open research question to determine the optimal balancing of $s$ and $r$.
@@ 5 Conclusion
In this work we have provided an analysis of the Krylov-aware low-rank approximation presented in [6]. In particular, we have shown that the Krylov-aware low-rank approximation to $f(\bm{A})$ is never worse than the approximation returned by the randomized SVD on $f(\bm{A})$, up to a polynomial approximation error factor. Furthermore, the numerical experiments demonstrate that the Krylov-aware low-rank approximation is significantly more accurate than the approximation returned by the randomized SVD on $f(\bm{A})$.
@@ Acknowledgements
This work was done while David Persson was a PhD-student at EPFL and supported by the SNSF research project Fast algorithms from low-rank updates, grant number: 200020_178806. This work is also contained in his PhD-thesis [34]. Tyler Chen and Christopher Musco were partially supported by NSF Award #2045590. The authors thank Daniel Kressner for helpful comments on this work.
@@
First note that by a standard Chebyshev interpolation bound [39, Lecture 20]
@@
Note that by the assumption on $s$ and a Stirling approximation $s!\geq\sqrt{2\pi s}\left(\frac{s}{e}\right)^{s}$ [37] we have $\frac{\gamma_{1,n}^{s}}{s!}<1$. Hence,
@@
Bounding $4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|\exp(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}$ is done identically to in Corollary 3.6. We proceed with bounding $\mathcal{E}(s;\exp(x))$. Define $\delta(x)=\frac{x-\lambda_{k+1}}{\lambda_{k+1}-\lambda_{n}}$. Define the polynomial $p(x)=(1+x-\lambda_{n})T_{s-2}\left(1+2\delta(x)\right)$, where $T_{s-2}$ is the Chebyshev polynomial of degree $s-2$. Hence, recalling the definition Eq. 16 of $\mathcal{E}(s;\exp(x))$, since $0\leq 1+x\leq\exp(x)$ for $x\geq 0$ and $|T_{s-2}(x)|\leq 1$ for $x\in[-1,1]$ we have,
@@
$\leq\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{T_{s-2}\left(1+2\delta(\lambda_{i})\right)}\right|^{2}.$ (22)
@@
Hence, for $\delta\geq 0$ we have $|T_{s-2}(1+2\delta)|\geq\frac{1}{2}(1+2\delta+2\sqrt{\delta+\delta^{2}})^{s-2}=:\frac{1}{2}h(\delta)^{s-2}$. Therefore,
$\max\limits_{i=1,\ldots,k}\left|\frac{e^{\lambda_{i}-\lambda_{n}}}{T_{s-2}\left(1+2\delta(\lambda_{i})\right)}\right|^{2}$ $\leq 4\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{h(\delta(\lambda_{i}))^{s-2}}\right|^{2}$
$=4\max\limits_{i=1,\ldots,k}\left(\exp(\lambda_{i}-\lambda_{n}-(s-2)\log(h(\delta(\lambda_{i}))))\right)^{2}.$
@@
$g(x):=x-\lambda_{n}-(s-2)\log(h(\delta(x)))$
is convex in $[\lambda_{k},\lambda_{1}]$ because $\log(h(x))$ is concave for $x\geq 0$ and $\delta(x)$ is linear positive function for any $x\geq\lambda_{k}$. Hence, the maximum of $g$ on the interval $[\lambda_{k},\lambda_{1}]$ is attained at either $x=\lambda_{k}$ or $x=\lambda_{1}$. The maximum of $g$ is attained at $x=\lambda_{k}$ whenever $g(\lambda_{1})\leq g(\lambda_{k})$ which happens whenever
$\frac{\lambda_{1}-\lambda_{k}}{\log\left(\frac{h(\delta(\lambda_{1}))}{h(\delta(\lambda_{k}))}\right)}+2\leq s.$ (23)
@@
$\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{T_{s-2}\left(1+2\delta(\lambda_{i})\right)}\right|^{2}\leq 4\frac{\exp(2(\lambda_{k}-\lambda_{n}))}{h(\delta(\lambda_{k}))^{2(s-2)}}.$ (24)
Now we use the argument from [29, p.21], which shows that $h(\delta)^{(s-2)}\geq 2^{\sqrt{\min\{1,2\delta\}}(s-2)-1}$. Plugging this inequality into Eq. 24 and then Eq. 22 and using the definition for $\gamma_{i,j}$ yields

### v2

@@ 5 Numerical experiments
In this section we compare the Krylov aware low-rank approximation (Algorithm 4), Algorithm 2 (assuming exact matvecs with $f(\bm{A})$) and Algorithm 3 (inexact matvecs with $f(\bm{A})$). All experiments have been performed in MATLAB (version 2024b) and scripts to reproduce the figures are available at https://github.com/davpersson/Krylov_aware_LRA.git.
@@ 5.1.1 Exponential integrator
The following example is taken from [persson_kressner_23]. Consider the following parabolic differential equation
[…]
@@
Hence, we can use Algorithm 4 to compute $\bm{Q}_{s}$ and $\bm{T}_{q}$ and use them to efficiently construct a rank $60$ approximation to $\exp(t\bm{A})$ for any $t$. Note that due to the well-known shift-invariance property of Krylov subspaces, as with most Krylov methods, an approximation can be obtained for many values of $t$ with little additional cost [Saad:1992, hochbruck_lubich, Hochbruck1998, Druskin1998].
@@ 5.1.2 Estrada index
For an (undirected) graph with adjacency matrix $\bm{A}$ the Estrada index is defined as $\operatorname{tr}(\exp(\bm{A}))$. It is used to measure the degree of protein folding [estrada]. One can estimate the Estrada index of a network by the Hutch++ algorithm or its variations [chen_hallman_23, epperly2023xtrace, hpp, ahpp], which requires computing a low-rank approximation of $\exp(\bm{A})$. Motivated by the numerical experiments in [hpp] we let $\bm{A}$ be the adjacency matrix of Roget’s Thesaurus semantic graph [roget].
@@
Figure 1: Comparing relative error (20) for the the approximations returned by Algorithm 4 without truncation (Krylov aware (untruncated)), Algorithm 4 with truncation back to rank $k$ (Krylov aware (truncated)), Algorithm 3 (randSVD), and Algorithm 2 (exact randSVD). The black line shows the optimal rank $k$ approximation relative Frobenius norm error. The rank parameter $k$ is visible as titles in the figures. In all experiments we set $\ell=k$.
@@ 5.1.3 Quantum spin system
We use an example from [epperly2023xtrace, Section 4.3], a similar example is found in [chen_hallman_23], in which we want to approximate $\exp(-\beta\bm{A})$ where
@@
Estimating the partition function $Z(\beta)=\operatorname{tr}(\exp(-\beta\bm{A}))$ is an important task in quantum mechanics [pfeuty1970one], which once again can benefit from computing a low-rank approximation of $\exp(-\beta\bm{A})$.
@@ 5.1.4 Synthetic example for the matrix logarithm
Figure 2: Comparing relative error (20) for the the approximations returned by Algorithm 4 without truncation (Krylov aware (untruncated)), Algorithm 4 with truncation back to rank $k$ (Krylov aware (truncated)), Algorithm 3 (randSVD), and Algorithm 2 (exact randSVD). The black line shows the optimal rank $k$ approximation relative Frobenius norm error. The rank parameter $k$ is visible as titles in the figures. In all experiments we set $\ell=k+5$.
5.2 Comparing relative errors of block methods
In this section we compare the error of the approximations returned by Algorithm 4, Algorithm 2, and Algorithm 3. If $\bm{C}$ is a low-rank approximation returned by one of the algorithms then we compare the relative error
@@ 5.2 Comparing relative errors of block methods
In all experiments we set the parameters in Algorithm 4 and Algorithm 3 to be $\ell=k$ or $\ell=k+5$ and $s=r$ so that the total number of matrix vector products with $\bm{A}$ is $2\ell s$. When we run Algorithm 2 we compute matvecs with $f(\bm{A})$ exactly, which cannot be done in practice. Hence, the results from this algorithm are only used as a reference for Algorithm 4 and Algorithm 3. The results are presented in Figure 1 for $\ell=k$ and Figure 2 for $\ell=k+5$. All results confirm that Algorithm 4 returns a more accurate approximation than Algorithm 3, and can even be more accurate than Algorithm 2. These results also confirm the findings of Chen & Hallman [chen_hallman_23], where it was shown that their Krylov aware version of Hutch++ yields more accurate approximations to $\operatorname{tr}(f(\bm{A}))$ compared to a naive implementation of Hutch++. As discussed in Section 1, the reason for the effectiveness of their method is due to a more accurate low-rank approximation.
5.3 Comparing relative errors of single vector methods
In this section we compare the relative error (20) of the approximations returned by Algorithm 4 and Algorithm 5; that is, we compare the block and single-vector version of the Krylov-aware low-rank approximation.
In Algorithm 4 we set the block-size to be $\ell=k$ and $s=r$. Consequently, Algorithm 4 performs at most $2sk$ matrix-vector products with $\bm{A}$ and returns the approximations $\textsf{{ALG}}(s,s;f)$ and $\textsf{{ALG}}_{k}(s,s;f)$. For comparison, we consider the output of Algorithm 5, namely $\widehat{\textsf{{ALG}}}((s-1)k,sk;f)$ and $\widehat{\textsf{{ALG}}}_{k}((s-1)k,sk;f)$, which also require at most $2sk$ matrix-vector products with $\bm{A}$. The results are presented in Figure 3. These result, shown in Figure 3, indicate that with the same matrix-vector products with $\bm{A}$, the single-vector version Algorithm 5 yields a more efficient approximation to $f(\bm{A})$ than Algorithm 4.
(a) Example from Section 5.1.1
(b) Example from Section 5.1.2
(c) Example from Section 5.1.3
(d) Example from Section 5.1.4
Figure 3: Comparing relative error (20) for the the approximations returned by Algorithm 4 without truncation (Krylov aware (untruncated)), Algorithm 4 with truncation back to rank $k$ (Krylov aware (truncated)), Algorithm 5 without truncation (SV Krylov aware (untruncated)), Algorithm 5 with truncation back to rank $k$ (SV Krylov aware (truncated)). The black line shows the optimal rank $k$ approximation relative Frobenius norm error. The rank parameter $k$ is visible as titles in the figures.
@@ 6 Conclusion
In this work we have provided an analysis of the Krylov-aware low-rank approximation presented in [chen_hallman_23]. In particular, we have shown that the Krylov-aware low-rank approximation to $f(\bm{A})$ is never worse than the approximation returned by the randomized SVD on $f(\bm{A})$, up to a polynomial approximation error factor, and we have presented an error analysis of the method. Numerical experiments further demonstrated that the Krylov-aware approach is often substantially more accurate than randomized SVD for approximating $f(\bm{A})$.
However, as highlighted in [thesis, Chapter 10], when using the same number of matvecs with $\bm{A}$ the approximation
$f(\bm{A})\approx\bm{Q}_{s}\llbracket f(\bm{T}_{s})\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}},\quad\bm{T}_{s}=\bm{Q}_{s}^{\textsf{{T}}}\bm{A}\bm{Q}_{s},$ (21)
where $\bm{Q}_{s}$ is an orthonormal basis for a Krylov subspace, can be even more accurate than the Krylov-aware low-rank approximation, at the same computational cost. This scheme corresponds to setting $r=0$ in Algorithm 4. It is also closely related to the method proposed in [persson_kressner_23, funnystrom2] for operator monotone functions.
While the analysis developed in this works does not yet provide theoretical justification for (21), extending the theory to cover (21) remains an important direction for future work. Nonetheless, our hope is that the present analysis serves as a stepping stone toward such a result.
@@ Acknowledgements
This work was done while David Persson was a PhD-student at EPFL and supported by the SNSF research project Fast algorithms from low-rank updates, grant number: 200020_178806. This work is also contained in his PhD-thesis [thesis]. Tyler Chen and Christopher Musco were partially supported by NSF Award #2045590. The authors thank Daniel Kressner for helpful comments on this work.
@@
First note that by a standard Chebyshev interpolation bound [stewart, Lecture 20]
@@
Note that by the assumption on $s$ and a Stirling approximation $s!\geq\sqrt{2\pi s}\left(\frac{s}{e}\right)^{s}$ [stirling] we have $\frac{\gamma_{1,n}^{s}}{s!}<1$. Hence,
@@
Bounding $4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|\exp(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}$ is done identically to in Corollary 3.6. We proceed with bounding $\mathcal{E}(s;\exp(x))$. Define ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(x)=\frac{x-\lambda_{k+1}}{\lambda_{k+1}-\lambda_{n}}$. Define the polynomial $p(x)=(1+x-\lambda_{n})T_{s-2}\left(1+2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(x)\right)$, where $T_{s-2}$ is the Chebyshev polynomial of degree s2s-2.999Polynomials similar to $p(x)$ are commonly used in the analysis of eigenvalue approximation using Krylov subspace methods; see [shmuel, paige1971computation, saad1980rates, underwood1975iterative]. Hence, recalling the definition Eq. 16 of $\mathcal{E}(s;\exp(x))$, since $0\leq 1+x\leq\exp(x)$ for $x\geq 0$ and $|T_{s-2}(x)|\leq 1$ for $x\in[-1,1]$ we have,
@@
$\leq\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{T_{s-2}\left(1+2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(\lambda_{i})\right)}\right|^{2}.$ (24)
@@
Hence, for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}\geq 0$ we have $|T_{s-2}(1+2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta})|\geq\frac{1}{2}(1+2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}+2\sqrt{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}^{2}})^{s-2}=:\frac{1}{2}h({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta})^{s-2}$. Therefore,
$\max\limits_{i=1,\ldots,k}\left|\frac{e^{\lambda_{i}-\lambda_{n}}}{T_{s-2}\left(1+2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(\lambda_{i})\right)}\right|^{2}$ $\leq 4\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{h({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(\lambda_{i}))^{s-2}}\right|^{2}$
$=4\max\limits_{i=1,\ldots,k}\left(\exp(\lambda_{i}-\lambda_{n}-(s-2)\log(h({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(\lambda_{i}))))\right)^{2}.$
@@
$g(x):=x-\lambda_{n}-(s-2)\log(h({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(x)))$
is convex in $\lambda_{k},\lambda_{1}]$ because $\log(h(x))$ is concave for $x\geq 0$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(x)$ is linear positive function for any $x\geq\lambda_{k}$. Hence, the maximum of $g$ on the interval $\lambda_{k},\lambda_{1}]$ is attained at either $x=\lambda_{k}$ or $x=\lambda_{1}$. The maximum of $g$ is attained at $x=\lambda_{k}$ whenever $g(\lambda_{1})\leq g(\lambda_{k})$ which happens whenever
$\frac{\lambda_{1}-\lambda_{k}}{\log\left(\frac{h({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(\lambda_{1}))}{h({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(\lambda_{k}))}\right)}+2\leq s.$ (25)
@@
$\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{T_{s-2}\left(1+2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(\lambda_{i})\right)}\right|^{2}\leq 4\frac{\exp(2(\lambda_{k}-\lambda_{n}))}{h({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}(\lambda_{k}))^{2(s-2)}}.$ (26)
Now we use the argument from [MM15, p.21], which shows that $h({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta})^{(s-2)}\geq 2^{\sqrt{\min\{1,2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta}\}}(s-2)-1}$. Plugging this inequality into Eq. 26 and then Eq. 24 and using the definition for $\gamma_{i,j}$ yields
@@
Noting $\exp(2\gamma_{k,n})\leq 2^{3\gamma_{k,n}}$ and inserting the definition of $\Gamma$ yields the desired result.
The second statement follows directly. Setting $\ell=2k+1$ and $s=O\left(\frac{\log(1/\varepsilon)+\gamma_{k,n}}{\sqrt{\Gamma}}\right)$ yields $2^{-2(s-2)\sqrt{\Gamma}+3\gamma_{k,n}}\frac{80k}{\ell-k-1}\leq\varepsilon$. Selecting $r$ as prescribed yields makes the polynomial approximation term $\leq\tau\|\exp(\bm{A})\|_{2}$ by a Stirling approximation on $(2r+2)!$ [stirling], yielding the desired result.
Appendix B Proof of Theorem 4.1
To prove Lemma B.1, we begin with the following simple lemma.
Lemma B.1.
Let $\bm{\omega}_{1}\in\mathbb{R}^{k}$ and $\bm{\omega}_{2}$ be standard Gaussian random vectors, and let $\bm{\omega}_{1,i}$ denote the $i^{\text{th}}$ entry of $\bm{\omega}_{1}$. For any fixed matrix $\bm{B}$, the following inequality holds with probability at least $1-\delta$
$\frac{\|\bm{B}\bm{\omega}_{2}\|_{2}^{2}}{\min\limits_{i=1,\ldots,k}\bm{\omega}_{1,i}^{2}}\leq\widehat{C}_{\delta}k^{2}\|\bm{B}\|_{\textsf{{F}}}^{2},$
where $\widehat{C}_{\delta}=\frac{2\pi(1+6\log(2/\delta))}{\delta^{2}}$ is as in Theorem 4.1.
Proof.
By [MeyerMuscoMusco:2024, Lemma 3.1] we have $\frac{1}{\min\limits_{i=1,\ldots,k}\bm{\omega}_{1,i}^{2}}\leq\frac{2\pi k^{2}}{\delta^{2}}$ with probability $1-\delta/2$. Furthermore, by a one-sided version of [cortinoviskressner, Theorem 1] we have $\mathbb{P}(\|\bm{B}\bm{\omega}_{2}\|_{2}^{2}\geq(1+x)\|\bm{B}\|_{\textsf{{F}}}^{2})\leq\exp\left(-\frac{x^{2}}{4(1+x)}\right)$. Setting $x=6\log(2/\delta)$ gives $\mathbb{P}(\|\bm{B}\bm{\omega}_{2}\|_{2}^{2}\geq(1+x)\|\bm{B}\|_{\textsf{{F}}}^{2})\leq\exp\left(-\frac{x^{2}}{4(1+x)}\right)\leq\delta/2$. Applying the union bound yields the desired result. ∎
With Lemma B.1 at hand, we are ready to prove Theorem 4.1.
Proof of Theorem 4.1.
Define
$\widehat{\bm{\Omega}}=\begin{bmatrix}\bm{\omega}&\bm{A}\bm{\omega}&\cdots&\bm{A}^{k-1}\bm{\omega}\end{bmatrix},\quad\widehat{\bm{\Omega}}_{k}=\bm{U}_{k}^{\textsf{{T}}}\widehat{\bm{\Omega}},\quad\widehat{\bm{\Omega}}_{n\setminus k}=\bm{U}_{n\setminus k}^{\textsf{{T}}}\widehat{\bm{\Omega}}$
By the discussion in Section 4 it suffices to show that for any polynomial $p\in\mathbb{P}_{s-1}$ we have
$\|p(\bm{\Lambda}_{n\setminus k})\widehat{\bm{\Omega}}_{n\setminus k}\widehat{\bm{\Omega}}_{k}^{-1}\|_{\textsf{{F}}}^{2}\leq\widehat{C}_{\delta}\frac{k^{4}}{\gamma_{\min}^{2(k-1)}}\|p(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}.$
Our proof follows the same idea in [MeyerMuscoMusco:2024, Theorem 3.1], but we take care to not introduce any dependencies on the matrix size $n$. We have
$\|p(\bm{\Lambda}_{n\setminus k})\widehat{\bm{\Omega}}_{n\setminus k}\widehat{\bm{\Omega}}_{k}^{-1}\|_{\textsf{{F}}}^{2}$ $\leq k\|p(\bm{\Lambda}_{n\setminus k})\widehat{\bm{\Omega}}_{n\setminus k}\widehat{\bm{\Omega}}_{k}^{-1}\|_{2}^{2}$
$=k\max\limits_{\bm{x}\neq\bm{0}}\frac{\|p(\bm{\Lambda}_{n\setminus k})\widehat{\bm{\Omega}}_{n\setminus k}\widehat{\bm{\Omega}}_{k}^{-1}\bm{x}\|_{2}^{2}}{\|\bm{x}\|_{2}^{2}}$
$=k\max\limits_{\bm{y}\neq\bm{0}}\frac{\|p(\bm{\Lambda}_{n\setminus k})\widehat{\bm{\Omega}}_{n\setminus k}\bm{y}\|_{2}^{2}}{\|\widehat{\bm{\Omega}}_{k}\bm{y}\|_{2}^{2}}$
$=k\max\limits_{g\in\mathbb{P}_{k-1},g\neq 0}\frac{\|p(\bm{\Lambda}_{n\setminus k})g(\bm{\Lambda}_{n\setminus k})\bm{\omega}_{n\setminus k}\|_{2}^{2}}{\|g(\bm{\Lambda}_{k})\bm{\omega}_{k}\|_{2}^{2}},$
where $\bm{\omega}_{k}=\bm{U}_{k}^{\textsf{{T}}}\bm{\omega}$ and $\bm{\omega}_{n\setminus k}=\bm{U}_{n\setminus k}^{\textsf{{T}}}\bm{\omega}$ are (independent) standard Gaussian random vectors. Using $p(\bm{\Lambda}_{n\setminus k})g(\bm{\Lambda}_{n\setminus k})=g(\bm{\Lambda}_{n\setminus k})p(\bm{\Lambda}_{n\setminus k})$, submultiplicativity of the Euclidean norm, and Lemma B.1 we have with probability $\geq 1-\delta$
$\|p(\bm{\Lambda}_{n\setminus k})\widehat{\bm{\Omega}}_{n\setminus k}\widehat{\bm{\Omega}}_{k}^{-1}\|_{\textsf{{F}}}^{2}$ $\leq k\max\limits_{g\in\mathbb{P}_{k-1},g\neq 0}\frac{\|p(\bm{\Lambda}_{n\setminus k})g(\bm{\Lambda}_{n\setminus k})\bm{\omega}_{n\setminus k}\|_{2}^{2}}{\|g(\bm{\Lambda}_{k})\bm{\omega}_{k}\|_{2}^{2}}$
$\leq k\frac{\|p(\bm{\Lambda}_{n\setminus k})\bm{\omega}_{n\setminus k}\|_{2}^{2}}{\min\limits_{i=1,\ldots,k}\bm{\omega}_{k,i}^{2}}\max\limits_{g\in\mathbb{P}_{k-1},g\neq 0}\frac{\|g(\bm{\Lambda}_{n\setminus k})\|_{2}^{2}}{\|g(\bm{\Lambda}_{k})\|_{\textsf{{F}}}^{2}}$
$\leq\widehat{C}_{\delta}k^{3}\max\limits_{g\in\mathbb{P}_{k-1},g\neq 0}\frac{\|g(\bm{\Lambda}_{n\setminus k})\|_{2}^{2}}{\|g(\bm{\Lambda}_{k})\|_{\textsf{{F}}}^{2}}.$ (27)
As done in the proof of [MeyerMuscoMusco:2024, Theorem 3.1], we write $g$ as an interpolating polynomial
$g(x)=\sum\limits_{i=1}^{k}g(\lambda_{i})\ell_{i}(x),\quad\ell_{i}(x)=\prod\limits_{t=1,t\neq i}^{k}\frac{x-\lambda_{t}}{\lambda_{i}-\lambda_{t}}.$
Note that a simple bound yields $|\ell_{i}(x)|\leq\frac{1}{\gamma_{\min}^{k-1}}$. Hence,
$\max\limits_{g\in\mathbb{P}_{k-1},g\neq 0}\frac{\|g(\bm{\Lambda}_{n\setminus k})\|_{2}^{2}}{\|g(\bm{\Lambda}_{k})\|_{\textsf{{F}}}^{2}}$ $=\max\limits_{g\in\mathbb{P}_{k-1},g\neq 0}\frac{\max\limits_{j=k+1,\ldots,n}\left|\sum\limits_{i=1}^{k}g(\lambda_{i})\ell_{i}(\lambda_{j})\right|^{2}}{\sum\limits_{i=1}^{k}g(\lambda_{i})^{2}}$
$\leq\max\limits_{j=k+1,\ldots,n}\left(\sum\limits_{i=1}^{k}|\ell_{i}(\lambda_{j})|\right)^{2}$
$\leq\max\limits_{x\in[\lambda_{\min},\lambda_{\max}]}\sum\limits_{i=1}^{k}|\ell_{i}(x)|^{2}$
$\leq\frac{k}{\gamma_{\min}^{2(k-1)}}.$ (28)
where we used the Cauchy-Schwarz inequality and that the bound for $|\ell_{i}(x)|$. Combining (28) with (27) yields the desired result. ∎